OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. let x be the larger number. The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (anonymous):

X+Y = 170 X-Y = 26 Solve.

OpenStudy (anonymous):

170/2 plus or minues 13 85 + 13 = 98 85 - 13 = 72 98+72 = 170 98-72 = 26